It was too embarassing to see that the actuation frequency was set at the violin mode frequency in order to avoid designing a new filter!?
Ooops, definitely my bad. I think I was the one who put in the PRM violin filter, so I should have recognized that frequency. However, I couldn't think of a reason why violin mode filters should be in the LSC filter banks, since we usually put them in C1:SUS-optic_LSC filter banks.
Anyhow, so that I don't make a mistake like that again, I was looking through all of the violin mode filters for all the optics, so I could write down the frequencies. The result: confusion.
Violin filters in C1:SUS-optic_LSC filter banks:
The PRM's violin mode filter is set correctly to 627.75Hz: elog 8533.
One of BS or SRM has probably been measured (presumably BS), since they have the same filters centered around 645Hz.
Neither ITM has a violin filter.
The ETMs have violin filters in the 440's, which I assume was correct back in the MOS days, before 2010.
Vio2 filters in C1:SUS-optic_LSC filter banks:
PRM, SRM, BS, ITMX, ITMY: Centered around 1285 Hz, which matches the violin notch frequencies in the BS and SRM.
ETMY: Centered around 883.5Hz, which matches the old 440Hz frequency
ETMX: Centered around 631Hz . So, this could have been measured, but it was put into the wrong filter module.
Koji tells me that we don't really need to worry about all these violin filters unless they are required (as with the PRM and the obnoxious hum a few weeks ago), so I 'm not going to do any measuring / adjusting of these filters for now.